of a rectangle whose sides measure 14 and 17?
Answer by jim_thompson5910(35256)

This rectangle has an area of 14*17 = 238


and it has a perimeter of 2*14 + 2*17 = 28 + 34 = 62


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim